# A Methodology for Incompleteness-Tolerant and Modular Gradual Semantics for Argumentative Statement Graphs

### v5

@@ 3.1 Methodology
$\bullet$ $\mathcal{A}^{1}=\{(\langle\Phi_{2},\Psi_{2}\rangle,x)\in\mathcal{A}(\alpha_{1})\times Prem(\alpha_{1})\>|\>\neg x\equiv\Psi_{2}\}$;
$\bullet$ $\mathcal{S}^{1}=\{(\langle\Phi_{2},\Psi_{2}\rangle,x)\in\mathcal{S}(\alpha_{1})\times Prem(\alpha_{1})\>|\>x\equiv\Psi_{2}\}$.
A statement’s premise graph contains its premises, attackers and supporters with the corresponding attacks and supports between them, as shown in Figure 2. Note that we do not specify $\tau^{1}$ at this point; we will do so in the specific instantiation of our methodology.
Our methodology, illustrated in Figure 2, is as follows.
@@
Figure 2: The conventional way (left) in which abstract GS ($\sigma_{*}$) are used, with a strength calculated for $\alpha_{1}$ using the strengths of its supporters ($\alpha_{2}$ and $\alpha_{4}$) and attacker ($\alpha_{3}$), and our approach (right), modular structured GS ($\sigma_{M}$), which calculate the strength, using $\sigma_{*}$, of each literal in the $\alpha_{1}$’s premises ($b$ and $c$) within the premise graph of $\alpha_{1}$, before $\sigma_{M}$ aggregates these strengths in a premise aggregation function to calculate the strength of $\alpha_{1}$.
The modular structured GS thus dialectically evaluates the evidence supporting or attacking each literal in a statement’s premise via an abstract GS using the premise graph, rather than based on the completeness of its support as in Jedwabny et al. (2020). These dialectical evaluations are then used to evaluate the statement with the premise aggregation function. This permits the incorporation of incomplete information, since the evidence for or against premises is combined dialectically starting from a statement’s weight, giving an approximation of the strength of a statement based on the available evidence, giving desirable behaviour in all three cases in Definition 9. This modular methodology permits instantiations tailored to individual applications based on choices in Definition 11. We now define one such instantiation, tailored to our restricted SGs of Definition 2, where premises are conjunctions, with its supporting abstract GS left unspecified.
@@
The DC semantics has been designed such that certain theoretical properties will be satisfied, as we will see in §3.2. In future work, modular structured GS could be instantiated for different SGs, e.g., for SGs with disjunctions as premises, a probabilistic sum could be used for the premise aggregation function, and for more complicated premises, a premise aggregation function could deal with each compound in turn.
@@ 3.2 Properties
We now discuss existing (from Jedwabny et al. (2020) [^7] Jedwabny et al. Jedwabny et al. (2020) also defined properties that are specific to CSTs, which we reformulate and assess in Appendix A.) or novel properties for GS in SGs. Note that the properties apply to either all or exclusively structured GS as specified.
The first property Jedwabny et al. (2020) is a fundamental requirement for GS, stating that a statement’s strength only depends on the statements that are connected to it.
@@
As in previous works Amgoud and Ben-Naim (2018); Amgoud et al. (2022), we would advocate that this property is satisfied since we would not expect information that is unrelated to a statement to have an effect on its strength.
Next Jedwabny et al. Jedwabny et al. (2020) require that the length of a proof in the logic should not affect the strength of a statement.
@@
$\bullet$ $\alpha_{1}=\langle\Phi_{1},\Psi_{1}\rangle$, where $\Phi_{1}\equiv x_{1}\wedge\ldots\wedge x_{n}\wedge x_{n+1}$ and $\Psi_{1}\equiv x$;
$\bullet$ $\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle$, where $\Phi_{2}\equiv x_{1}\wedge\ldots\wedge x_{n}$ and $\Psi_{2}\equiv x^{\prime}$;
$\bullet$ $\alpha_{3}=\langle\Phi_{3},\Psi_{3}\rangle$, where $\Phi_{3}\equiv x^{\prime}\wedge x_{n+1}$ and $\Psi_{3}\equiv x$;
$\bullet$ $\nexists\langle\Phi_{4},\Psi_{4}\rangle$ where $x^{\prime}\in Prem(\langle\Phi_{4},\Psi_{4}\rangle)$ or $x^{\prime}\equiv\Psi_{4}$;
$\bullet$ $\tau(\alpha_{1})=\tau(\alpha_{3})$ and $\tau(\alpha_{2})=1$;
@@
This property is indeed intuitive in a number of settings where we would prioritise the logical consistency over approximate evaluations of statements. However, in some settings, e.g., when modelling human discourse, it may be the case that the same logical conclusions, when reached over the course of multiple statements, are weaker, or even stronger, than those stated succinctly (e.g., see the study of priority in Yin et al. (2024)). For example, given the statements instantiated in rewriting, consider the effect of adding an statement $\alpha_{4}=\langle\Phi_{4},\Psi_{4}\rangle$, where $\Psi\equiv\neg x_{1}$. In some settings may be desirable that the weakening effect of $\alpha_{4}$ on $\alpha_{3}$ is lesser than that on $\alpha_{1}$ due to the distance. We leave a formalisation of this (non-trivial) property to future work. We would, however, argue that rewriting may not be universally applicable.
@@
Next, Jedwabny et al. Jedwabny et al. (2020) state that a statement without at least one supporter for each of its premises is rejected.
@@
We now propose more novel properties that offer alternatives to provability, firstly considering the “initial state” of statements, i.e., that the strength of a statement with no attackers or supporters should be equal to its weight, as in properties defined for abstract GS, namely stability Amgoud and Ben-Naim (2018) and balance Baroni et al. (2019). It is also related to logical void precedence Heyninck et al. (2023).
@@
First, we consider the effect of adding a statement that has been assigned a zero strength, i.e., that it should not affect the strength of any statement it attacks or supports, in line with neutrality Amgoud and Ben-Naim (2018) for abstract GS.
@@
This property, as with stability, takes a notably different approach to the T-norm semantics given their focus on CSTs, namely that the strengths of statements are calculated recursively, but should be unaffected by rejected statements.
@@
Let us revisit our running example and let $\alpha_{7}\!=\!\langle\top,d\rangle$ (“It is a fact that CO2 measurements are unreliable”) with $\tau(\alpha_{7})\!=\!0$. Then, under the DC and T-norm-p semantics, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{7})\!=\!\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{7})\!=\!0$, i.e., $\alpha_{7}$ is rejected. However, the strength of $\alpha_{4}$ under DC is $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{4})\!=\!0.8$ (unchanged). However, under the T-norm-p semantics, it is still $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{4})\!=\!0$ (rejected). Under the DC semantics, $\alpha_{7}$ is rejected and, thus, has no influence on the strength of $\alpha_{4}$. In contrast, the rejection of $\alpha_{7}$ under T-norm-p leads to a rejection of $\alpha_{4}$ due to the multiplicative property, thus violating neutrality.
The next two properties consider the effect of an attacking or supporting statement being added which has not been rejected, taking inspiration from bi-variate monotony Amgoud and Ben-Naim (2018), monotonicity Baroni et al. (2019) and P3 in Skiba et al. (2023). The first requires that adding an attacker should not strengthen a statement.
@@
Next, we consider the effect of increasing the strengths of statements that attack or support other statements, as in bi-variate reinforcement Amgoud and Ben-Naim (2018) and monotonicity Baroni et al. (2019). First, strengthening an attacker should not strengthen a statement.
@@
These properties assess added statements’ effects, conditioning on their strengths rather than weights and concerning direct attackers and supporters rather than CSTs, thus accommodating statements with all three levels of completeness.
@@
Next, a rejected premise causes a statement to be rejected.
@@
A structured GS $\sigma$ satisfies rejected premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\exists x\in Prem(\alpha_{1})$ such that $\exists\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{A}(\alpha_{1})$ with $\Psi_{2}\equiv\neg x$ and $\sigma^{\mathcal{G}}(\alpha_{2})=1$ and $\nexists\alpha_{3}=\langle\Phi_{3},\Psi_{3}\rangle\in\mathcal{S}(\alpha_{1})$ with $\Psi_{3}\equiv x$ and $\sigma^{\mathcal{G}}(\alpha_{3})>0$, $\sigma^{\mathcal{G}}(\alpha_{1})=0$.
For partially-complete or incomplete statements especially, this property provides an alternative reason for requiring rejection, i.e., there is maximal evidence to the contrary of one of its premises since they are, critically, part of a conjunction, as opposed to provability, where this is the case unless every premise in the statement is supported. Note that rejected premise is also desirable for complete statements.
@@
In our running example, let $\alpha_{8}=\langle e\wedge f,\neg b\rangle$ (“Extreme weather events are increasing and solar activity is high, thus global temperatures are not increasing”) with $\tau(\alpha_{8})\!=\!0.5$. Now, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{8})=0.707$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{8})=0$. Let us now add a new statement $\alpha_{9}=\langle\top,\neg f\rangle$ (“It is a fact that solar activity is not high”) with $\tau(\alpha_{9})=1$. Then, the strength of $\alpha_{8}$ is $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{7})=\sigma^{\mathcal{G}^{1}}_{D}(e)\times\sigma^{\mathcal{G}^{1}}_{D}(g)=0.707\times 0=0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{8})=0$. Both semantics now reject $\alpha_{8}$, but for different reasons. The DC semantics rejects it because one of its premises ($f$) is fully contradicted by $\alpha_{9}$, while the T-norm-p semantics already rejected it due to incomplete support.
Next, statements with accepted premises should be also.
@@
A structured GS $\sigma$ satisfies accepted premises iff for any $\alpha_{1}\in\mathcal{X}$, where $\forall x\in Prem(\alpha_{1})$, $\nexists\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{A}(\alpha_{1})$ such that $\Psi_{2}\!\!\equiv\!\!\neg x$ and $\sigma^{\mathcal{G}}(\alpha_{2})\!\!>\!\!0$ and $\exists\alpha_{3}\!=\!\langle\Phi_{3},\Psi_{3}\rangle\!\in\!\mathcal{S}(\alpha_{1})$ such that $\Psi_{3}\!\equiv\!x$ and $\sigma^{\mathcal{G}}(\alpha_{3})\!=\!1$, $\sigma^{\mathcal{G}}(\alpha_{1})\!=\!1$.
@@
Finally, attackers and supporters may affect statements symmetrically, i.e., two statements’ premises that negate one another have “opposite” strengths, as in duality Potyka (2018) and franklin Amgoud and Ben-Naim (2018). [^8] For the DC semantics, mirroring’s conditions are triggered only for statements with single literals in their premises. However, this may not be the case for other modular structured GS.
@@
In our running example, let $\alpha_{10}=\langle\neg a,\neg e\rangle$ (“CO2 emissions are not rising, thus extreme weather events are not increasing”) with $\tau(\alpha_{5})=\tau(\alpha_{10})\!=\!0.5$. Then, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{10})\!=\!1-\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{5})\!=\!1-0.55\!=\!0.45$, and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{5})\!=\!0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{10})\!=\!0.45$. Despite $\alpha_{5}$ and $\alpha_{10}$ having premises that negate one another and equal base weights, their strengths are not complementary (do not sum to 1) under the T-norm-p semantics. In contrast, the DC semantics show a desirable duality in the effects of the attackers and supporters.

### v6

@@ 4.1 Methodology
  • $\mathcal{A}^{1}=\{(\langle\Phi_{2},\Psi_{2}\rangle,x)\in\mathcal{A}(\alpha_{1})\times Prem(\alpha_{1})\>|\>\neg x\equiv\Psi_{2}\}$;
  • $\mathcal{S}^{1}=\{(\langle\Phi_{2},\Psi_{2}\rangle,x)\in\mathcal{S}(\alpha_{1})\times Prem(\alpha_{1})\>|\>x\equiv\Psi_{2}\}$.
A statement’s premise graph contains a new set of statements representing the statement’s premises, attackers and supporters from the original SG, as shown in the dashed blue box on the right of Figure 3. Note that we do not specify $\tau^{1}$ at this point; we will do so in the specific instantiation of our methodology. In introducing a premise graph, we allow for the literals in statements’ premises to be evaluated by abstract gradual semantics, a core tenet of our approach.
Our methodology, illustrated in Figure 3, is as follows. [^8] Given the modular nature of our approach, we leave a study of the complexities of the resulting GS to future work.
@@
Figure 3: The conventional way (left) in which abstract GS ($\sigma_{*}$) are used for SGs (centre), with a strength calculated for $\alpha_{1}$ using the strengths of its supporters ($\alpha_{2}$ and $\alpha_{4}$) and attacker ($\alpha_{3}$), and our approach (right), modular structured GS ($\sigma_{M}$), which calculate the strength, using $\sigma_{*}$, of each literal in $\alpha_{1}$’s premises ($b$ and $c$) within the premise graph of $\alpha_{1}$, before $\sigma_{M}$ aggregates these strengths in a premise aggregation function to calculate the strength of $\alpha_{1}$.
The modular structured GS thus dialectically evaluates the evidence supporting or attacking each literal in a statement’s premise via an abstract GS using the premise graph, rather than based on the completeness of its support as in (?). These dialectical evaluations are then used to evaluate the statement with the premise aggregation function. This permits the incorporation of incomplete information, since the evidence for or against premises is combined dialectically starting from a statement’s weight, giving an approximation of the strength of a statement based on the available evidence, giving desirable behaviour in all three cases in Definition 9. This modular methodology permits instantiations tailored to individual applications based on choices in Definition 11. We now define one such instantiation, tailored to our restricted SGs of Definition 2, where premises are conjunctions, with its supporting abstract GS left unspecified.
@@
Note that conjunctive premises represent a challenge for handling incomplete information, as the failure of any single premise invalidates the entire statement. For that matter, the DC semantics has been designed such that certain desirable theoretical properties will be satisfied, as we will see in §4.2. Importantly, the DC semantics represents a natural starting point for our methodology because conjunctive premises are typically abound in real-world argumentative reasoning. For instance, when people construct arguments, they typically combine multiple pieces of evidence conjunctively to support their claims. This pattern is evident in scientific argumentation, legal reasoning and everyday discourse, where people naturally aggregate supporting facts with “and” connectives (?). In future work, modular structured GS could be instantiated for different structures of SGs, e.g., for SGs with disjunctions as premises, a probabilistic sum could be used for the premise aggregation function. This would align with the intuition that multiple independent supporting lines of reasoning may strengthen a statement’s overall strength. For other logical structures such as statements in conjunctive normal form, a hierarchical aggregation approach could be employed as explored, i.e., first combining literals within clauses using disjunctive operators, then aggregating clauses conjunctively.
@@ 4.2 Properties
We now discuss existing (from (?) [^9] Jedwabny et al. (?) also defined properties that are specific to CSTs, which we reformulate and assess in Appendix A.) and novel properties for GS in SGs. Note that the properties apply to either all or exclusively structured GS as specified.
The first property (?) is a fundamental requirement for GS, stating that a statement’s strength only depends on the statements that are connected to it.
@@
As in previous works (??), we would advocate that this property is satisfied in the vast majority of settings since we would not expect information that is unrelated to a statement to have an effect on its strength.
Next, Jedwabny et al. (?) introduced a property which requires that the length of a proof in the logic should not affect the strength of a statement.
@@
  • $\alpha_{1}=\langle\Phi_{1},\Psi_{1}\rangle$, where $\Phi_{1}\equiv x_{1}\wedge\ldots\wedge x_{n}\wedge x_{n+1}$ and $\Psi_{1}\equiv x$;
  • $\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle$, where $\Phi_{2}\equiv x_{1}\wedge\ldots\wedge x_{n}$ and $\Psi_{2}\equiv x^{\prime}$;
  • $\alpha_{3}=\langle\Phi_{3},\Psi_{3}\rangle$, where $\Phi_{3}\equiv x^{\prime}\wedge x_{n+1}$ and $\Psi_{3}\equiv x$;
  • $\nexists\langle\Phi_{4},\Psi_{4}\rangle$ where $x^{\prime}\in Prem(\langle\Phi_{4},\Psi_{4}\rangle)$ or $x^{\prime}\equiv\Psi_{4}$;
  • $\tau(\alpha_{1})=\tau(\alpha_{3})$ and $\tau(\alpha_{2})=1$;
@@
This property is indeed intuitive in a number of settings where we would prioritise the logical consistency over approximate evaluations of statements. However, in some settings, e.g., when modelling human discourse, it may be the case that the same logical conclusions, when reached over the course of multiple statements, are weaker, or even stronger, than those stated succinctly (e.g., see the study of priority in (?)). For example, given the statements instantiated in rewriting, consider the effect of adding an statement $\alpha_{4}=\langle\Phi_{4},\Psi_{4}\rangle$, where $\Psi\equiv\neg x_{1}$. In some settings may be desirable that the weakening effect of $\alpha_{4}$ on $\alpha_{3}$ is lesser than that on $\alpha_{1}$ due to the distance. We leave a formalisation of this (non-trivial) property to future work. We would, however, argue that rewriting may not be universally applicable.
@@
Next, Jedwabny et al. (?) state that a statement without at least one supporter for each of its premises is bottom-strength.
@@
We now propose more novel properties that offer alternatives to provability, firstly considering the “initial state” of statements, i.e., that the strength of a statement with no attackers or supporters should be equal to its weight, as in properties defined for abstract GS, namely stability (?) and balance (?). It is also related to logical void precedence (?).
@@
First, we consider the effect of adding a statement that has been assigned a zero strength, i.e., that it should not affect the strength of any statement it attacks or supports, in line with neutrality (?) for abstract GS.
@@
This property, as with stability, takes a notably different approach to the T-norm semantics given their focus on CSTs, namely that the strengths of statements are calculated recursively, but should be unaffected by bottom-strength statements.
@@
Let us revisit our running example and let $\alpha_{7}\!=\!\langle\top,d\rangle$ (“It is a fact that CO2 measurements are unreliable”) with $\tau(\alpha_{7})\!=\!0$. Then, under the DC and T-norm-p semantics, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{7})\!=\!\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{7})\!=\!0$, i.e., $\alpha_{7}$ is bottom-strength. However, the strength of $\alpha_{4}$ under DC is $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{4})\!=\!0.8$ (unchanged). However, under the T-norm-p semantics, it is still $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{4})\!=\!0$ (bottom-strength). Under the DC semantics, $\alpha_{7}$ is bottom-strength and, thus, has no influence on the strength of $\alpha_{4}$. In contrast, the bottom-strength of $\alpha_{7}$ under T-norm-p leads to the bottom-strength of $\alpha_{4}$ due to the multiplicative property, thus violating neutrality.
The next two properties consider the effect of an attacking or supporting statement being added which is not bottom-strength, taking inspiration from bi-variate monotony (?), monotonicity (?) and P3 in (?). The first requires that adding an attacker should not strengthen a statement.
@@
Next, we consider the effect of increasing the strengths of statements that attack or support other statements, as in bi-variate reinforcement (?) and monotonicity (?). First, strengthening an attacker should not strengthen a statement.
@@
These properties assess added statements’ effects, conditioning on their strengths rather than weights and direct attackers and supporters rather than CSTs, thus accommodating statements of all three levels of completeness.
@@
Next, a bottom-strength premise causes a statement to be bottom-strength. This property is related to argument death (?).
@@
A structured GS $\sigma$ satisfies bottom-strength premise iff for any $\alpha_{1}\in\mathcal{X}$, where $\exists x\in Prem(\alpha_{1})$ such that $\exists\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{A}(\alpha_{1})$ with $\Psi_{2}\equiv\neg x$ and $\sigma^{\mathcal{G}}(\alpha_{2})=1$ and $\nexists\alpha_{3}=\langle\Phi_{3},\Psi_{3}\rangle\in\mathcal{S}(\alpha_{1})$ with $\Psi_{3}\equiv x$ and $\sigma^{\mathcal{G}}(\alpha_{3})>0$, $\sigma^{\mathcal{G}}(\alpha_{1})=0$.
For partially-complete or incomplete statements especially, this property provides an alternative reason for requiring rejection, i.e., there is maximal evidence to the contrary of one of its premises since they are, critically, part of a conjunction, as opposed to provability, where this is the case unless every premise in the statement is supported. Note that bottom-strength premise is also desirable for complete statements.
@@
In our running example, let $\alpha_{8}=\langle e\wedge f,\neg b\rangle$ (“Extreme weather events are increasing and solar activity is high, thus global temperatures are not increasing”) with $\tau(\alpha_{8})\!=\!0.5$. Now, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{8})=0.707$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{8})=0$. Let us now add a new statement $\alpha_{9}=\langle\top,\neg f\rangle$ (“It is a fact that solar activity is not high”) with $\tau(\alpha_{9})=1$. Then, the strength of $\alpha_{8}$ is $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{7})=\sigma^{\mathcal{G}^{1}}_{D}(e)\times\sigma^{\mathcal{G}^{1}}_{D}(g)=0.707\times 0=0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{8})=0$. Both semantics now reject $\alpha_{8}$, but for different reasons. The DC semantics rejects it because one of its premises ($f$) is fully contradicted by $\alpha_{9}$, while the T-norm-p semantics already gives it bottom-strength due to incomplete support.
Next, statements with universally top-strength premises should also be top-strength.
@@
A structured GS $\sigma$ satisfies top-strength premises iff for any $\alpha_{1}\in\mathcal{X}$, where $\forall x\in Prem(\alpha_{1})$, $\nexists\alpha_{2}=\langle\Phi_{2},\Psi_{2}\rangle\in\mathcal{A}(\alpha_{1})$ such that $\Psi_{2}\!\!\equiv\!\!\neg x$ and $\sigma^{\mathcal{G}}(\alpha_{2})\!\!>\!\!0$ and $\exists\alpha_{3}\!=\!\langle\Phi_{3},\Psi_{3}\rangle\!\in\!\mathcal{S}(\alpha_{1})$ such that $\Psi_{3}\!\equiv\!x$ and $\sigma^{\mathcal{G}}(\alpha_{3})\!=\!1$, $\sigma^{\mathcal{G}}(\alpha_{1})\!=\!1$.
@@
Finally, attackers and supporters may affect statements symmetrically, i.e., two statements’ premises that negate one another have “opposite” strengths, as in duality (?) and franklin (?). [^10] For the DC semantics, mirroring’s conditions are triggered only for statements with single literals in their premises. However, this may not be the case for other modular structured GS.
@@
In our running example, let $\alpha_{10}=\langle\neg a,\neg e\rangle$ (“CO2 emissions are not rising, thus extreme weather events are not increasing”) with $\tau(\alpha_{5})\!=\!\tau(\alpha_{10})\!=\!0.5$. Then, $\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{10})\!=\!1-\sigma^{\mathcal{G}}_{\wedge_{D}}(\alpha_{5})\!=\!1-0.55\!=\!0.45$, and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{5})\!=\!0$ and $\sigma^{\mathcal{G}}_{T_{p}}(\alpha_{10})\!=\!0.45$. Despite $\alpha_{5}$ and $\alpha_{10}$ having premises that negate one another and equal base weights, their strengths are not complementary (do not sum to 1) under T-norm-p semantics. In contrast, the DC semantics show a desirable duality in attackers’ and supporters’ effects.